OpenStudy (anonymous):

Part A: Wilson bought fruit that weighs 2 and 1 over 2 pounds. How many ounces does the fruit weigh? Show your work. (5 points) [16 ounces = 1 pound] Part B: A running tap dispenses 0.14 gallons of water every second. How many pints of water is dispensed after 50 seconds? Show your work. (5 points) [1 gallon = 4 quarts, 1 quart = 2 pints]

OpenStudy (anonymous):

@llamb25

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Pandabear78963

OpenStudy (anonymous):

@HaileyD @hybrik @hhopke @hk420 @happilyfalling @hype.child @HackberryAbby18 @kelseymariah @kpw1 @kattck @Kathybrock @Keigh2015 @kenialove @@*ImaBeMe* @Legends @Loonyone @lovejenniferv @lovejenniferv @lily1021 @LynFran @Lynienicole @ploopgraal @pistol5002 @poodahhh @Pandabear78963 @psychoshadow_kat @paige0123 @@LiveLaughDie @LivingK1ng @livlovhope @okodksokds @OpenStudy_Police @onewiththemoon @oOMadisonOo @obeyme913 @OGthug @oklauren @okodksokds

OpenStudy (lynfran):

if 16 ounce = 1 pound, then 16 ounce/ 2 = 1 over 2 pounds

OpenStudy (lynfran):

what 16/2=?

OpenStudy (anonymous):

8

OpenStudy (lynfran):

so 8 ounces =1/2 pounds

OpenStudy (lynfran):

no we add 16 ounce + 16 ounce + 8 ounce = 2 and 1 over 2 pounds

OpenStudy (anonymous):

thx @LynFran

OpenStudy (lynfran):

ok for part B.. since the running tap dispenses 0.14 gallon per second ,then in 50 seconds it will dispense 0.14* 50= 7 gallon ...now we convert using the above ... since 1 gallon =4 quarts , then 7 gallon = 7*4 quarts

OpenStudy (lynfran):

what 7*4=?

OpenStudy (anonymous):

28 sorry

OpenStudy (lynfran):

so 7 gallon= 28 quarts and since 1 quart=2 pints.., then 28 quarts= 28*2 pint

OpenStudy (lynfran):

what 28*2 =?

OpenStudy (anonymous):

56

OpenStudy (lynfran):

yep that shd be it 56 pint of water is dispense

OpenStudy (anonymous):

thx u sooooooo much

OpenStudy (lynfran):

welcome